# Gradient Aligned Regression via Pairwise Losses

### v3

@@ 1 Introduction
As one of the most fundamental tasks in Machine Learning (ML) field, regression stands out as a powerful tool for understanding and modeling the relationships from complex data features to continuous data labels. Regression techniques have been widely utilized in many areas, such as computer vision (Moschoglou et al., 2017; Niu et al., 2016), drug discovery (Dara et al., 2022), economics and finances (Benkraiem and Zopounidis, 2021), environmental forecast (Zhu et al., 2018), material science (Stanev et al., 2018). For centuries, Mean Squared Error (MSE) and Mean Absolute Error (MAE) have been adopted to measure the distance from the ML model predictions to the ground truths. Consequently, researchers have naturally proposed to optimize the least squared error and least absolute error (Legendre, 1806; Dodge, 2008). Since the last century, to improve model stability and control model complexity, researchers have proposed various model regularizations based on MSE loss, resulting in multiple variants such as Ridge Regression, LASSO, and Elastic Net (Tikhonov et al., 1943; Tibshirani, 1996; Zou and Hastie, 2005). Along this line, the Huber loss was proposed to trade off between MAE and MSE, aiming to mitigate the issue of MSE being sensitive to outliers (Huber, 1992). All these conventional loss functions and many of their variants focus on minimizing the difference between model prediction and ground truth for each individual data sample, but do not explicitly model and capture the relationships for multiple data samples.
@@ 3 Method
The conventional regression task optimizes the following objective:
@@
$\mathcal{L}^{\text{MAE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}|\sigma^{f}_{\mathbf{x}_{i}}|,\quad L^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}})^{2}.$ (2)
In this work, we will focus on regression on the clean data setting without considering noises, outliers or distributional shifts, etc.
@@ 3.1 Motivations
Motivated by the second point, we explicitly enforce the pairwise prediction difference $f(\mathbf{x}_{i})-f(\mathbf{x}_{j})$ to be close to the pairwise label difference $y_{i}-y_{j}$ through two pairwise losses in the label space: the first directly matches the original label difference, while the second is a relaxed version with a scaling factor. Next, we prove that the proposed two pairwise losses are equivalent to the variance of prediction errors and the negative Pearson correlation coefficient between predictions and ground truths, which addresses the concern in the first point. Moreover, the equivalent forms reduce the quadratic time complexity for the proposed pairwise losses to linear complexity. Last but not least, we provide theoretical insight between learning the pairwise label difference and learning the gradients of the ground truth function.
@@ 3.2 Pairwise Losses
When choose individual loss $\ell$ as $\frac{1}{2}$ squared error function:
@@
$\mathcal{L}^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}}+\bar{\sigma}^{f}_{\mathbf{x}})^{2}=\frac{1}{N}\sum_{i=1}^{N}[(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}})^{2}+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}]=\text{Var}(\sigma^{f}_{\mathbf{x}})+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}.$
@@
Remark: The proof mainly depends on Mean Value Theorem, Limits and L’Hôpital’s rule. Besides of fitting the prediction $f(\mathbf{x}_{i})$ to be close to ground truth $y_{i}$, i.e. $f(\mathbf{x})\approx\mathcal{Y}(\mathbf{x}),~{}\forall(\mathbf{x},\mathcal{Y}(\mathbf{x}))\in\mathcal{D}$, we additionally make the model to explicitly capture the gradient (and the higher order gradients) of the ground truth function, i.e. $\nabla^{k}f(\mathbf{x})\approx\nabla^{k}\mathcal{Y}(\mathbf{x}),~{}k=\{1,2,...\}$, with the proposed pairwise losses.
@@ 4.1 Synthetic Experiments
Experimental Settings: we include the full details for experimental settings in the Appendix B.3.
@@
Table 1: Testing performance on the tabular datasets. Mean and standard deviation are reported. ‘Gains’ includes the percentages that GAR outperform the MAE and the best competitor (in parenthesis). The ‘p-values’ are calculated based on Student’s t-test for the ‘Gains’. p-value < 0.05 usually means significant difference. $\downarrow$ means the smaller the better; $\uparrow$ means the larger the better.
@@ 4.2 Benchmark Experiments
Experimental Settings: we follow the standard procedure for data splitting, cross validation and hyper-parameter tuning. Due to limited space, we include the full details in Appendix B.7.
@@ 4.3 Further Analysis
We additionally provide practical running time comparison, ablation studies on different components and sensitivity analysis on $\alpha$ for GAR in this subsection.
Running Time Comparison: we run all compared methods sequentially on an exclusive cluster node with AMD EPYC 7402 24-Core Processor 2.0 GHz. Each method runs 1500 epochs for the five tasks on the tabular datasets. The results are summarized in Table 2. We find that GAR is as efficient as the most efficient conventional regression loss, such as MAE, MSE, etc. It is consistent with the theoretical result that GAR only takes linear time complexity. On the other hand, the pairwise latent feature based regression methods, such as RankSim, ConR and RNC, are clearly slower than GAR.
Table 2: Running time (seconds/epoch) on five benchmark datasets. Repeated 1500 epochs, mean and standard deviation are reported. ‘CCS’: Concrete Compressive Strength.
Dataset CCS Wine Quality Parkinson (Total) Super Conductivity IC50
MAE 0.093(0.075) 0.209(0.221) 0.303(0.412) 0.604(0.549) 0.126(0.09)
MSE 0.095(0.087) 0.23(0.249) 0.296(0.374) 0.959(0.653) 0.103(0.074)
Huber 0.088(0.067) 0.226(0.254) 0.303(0.396) 0.646(0.636) 0.13(0.093)
Focal (MAE) 0.091(0.067) 0.224(0.252) 0.292(0.374) 0.627(0.503) 0.134(0.101)
Focal (MSE) 0.087(0.067) 0.221(0.248) 0.302(0.383) 0.859(0.53) 0.129(0.088)
RankSim 0.315(0.296) 0.244(0.264) 1.67(0.981) 6.273(2.883) 0.358(0.28)
RNC 0.864(0.236) 3.275(1.16) 3.308(1.034) 10.342(2.296) 0.78(0.469)
ConR 0.165(0.253) 0.401(0.647) 0.445(0.68) 1.053(0.916) 0.154(0.119)
GAR (ours) 0.080(0.058) 0.223(0.251) 0.278(0.34) 0.613(0.512) 0.127(0.092)
@@
Sensitivity on $\alpha$: we conduct extensive hyper-parameter search for $\alpha$ in {0.1, 0.2, 0.25, 0.4, 0.5, 0.8, 1, 1.25, 2, 2.5, 4, 5, 10}. The initial learning rate and weight decay are still searched in the same previous sets that are described in Appendix B.7. We include the full results in Appendix B.10. From the box-plot, GAR is not very sensitive to $\alpha$ but it is still suggested to tune it in the [0.1, 10] range.
@@ 5 Limitations
The scope for this work is limited to improving learning label pairwise difference for regression on clean data setting, without considering noises, outliers or distributional shifts. Based on the Theorem 4, we relate GAR to capture the gradient of the ground truth function. However, GAR might be misled if the training data doesn’t reflect the underlying ground truth function. It would be an interesting future work to research the robustness for GAR under extensive settings.
@@ B.6 Backbone Model Complexity Variation on Sine Dataset
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ B.7 Experimental Setting on Real World Datasets
For tabular datasets, we uniformly randomly split 20% data as testing; the remaining 80% as training and validation, where we conduct 5-fold-cross-validation with random seed set as 123. The total training epochs is set as 100 and batch size is set as 256. The weight decay for each method is tuned in {1e-3, 1e-4, 1e-5}; we utilize SGD with momentum (set as 0.9) optimizer and tune the initial learning rate for baseline method in {1e-1, 1e-2, 1e-3, 1e-4, 1e-5}, which is stage-wised decreased by 10 folds at the end of 50-th and 75-th epoch. The switching hyper-parameter $\delta$ for Huber loss and the scaling hyper-parameter $\beta$ for Focal (MAE) or Focal (MSE) loss are tuned in {0.25, 1, 4}. The interpolation hyper-parameter $\lambda$ for RankSim is tuned in {0.5, 1, 2} and the balancing hyper-parameter $\gamma$ is fixed as 100 as suggested by their sensitivity study in their Appendix C.4 (Gong et al., 2022). The temperature hyper-parameter for RNC is tuned in {1,2,4}; the first 50 epochs are used for RNC pre-training and the remaining 50 epochs are used for fine-tuning with MAE loss. The linear combination hyper-parameters $\alpha$ is fixed as 1, $\beta$ is tuned in {0.2, 1, 4} for ConR, as suggested by the ablation studies in their Appendix A.5 (Keramati et al., 2023). The robust reconciliation hyper-parameter $\alpha$ for GAR is tuned in {0.1, 1, 10}. The model performance is evaluated with the following 4 metrics: MAE, RMSE, Pearson correlation coefficient, Spearman’s rank correlation coefficient. The best testing performance for each baseline method on each training fold is decided by its best validation performance on the same training fold over (initial learning rate, weight decay, method-special-hyper-parameter, the epoch for evaluation) combinations. For IC50 with 15 targets, we report the averaged value across all targets. For AgeDB dataset, we follow the same setting for RNC linear probe task and similar setting for training from scratch task. We include the complete details in Appendix B.8. Because RNC is a pre-training method, it doesn’t have evaluation on training from scratch task. Instead, all baselines in linear probe are based on RNC pre-trained latent features. Because RankSim and ConR are proposed to manipulate latent feature, they don’t have evaluation on linear probe task where the trainable model is linear. We additionally include Coefficient of determination (R2) as an evaluation metric for AgeDB besides of the previous 4 evaluation metrics, so as to compare with prior researches.

### v4

@@ 1 Introduction
As one of the most fundamental tasks in Machine Learning (ML) field, regression stands out as a powerful tool for understanding and modeling the relationships from complex data features to continuous data labels. Regression techniques have been widely utilized in many areas, such as computer vision (Moschoglou et al., 2017; Niu et al., 2016), drug discovery (Dara et al., 2022), economics and finances (Benkraiem and Zopounidis, 2021), environmental forecast (Zhu et al., 2018), material science (Stanev et al., 2018). For centuries, Mean Squared Error (MSE) and Mean Absolute Error (MAE) have been adopted to measure the distance from the ML model predictions to the ground truths. Consequently, researchers have naturally proposed to optimize the least squared error and least absolute error (Legendre, 1806; Dodge, 2008). Since the last century, to improve model stability and control model complexity, researchers have proposed various model regularizations based on MSE loss, resulting in multiple variants such as Ridge Regression, LASSO, and Elastic Net (Tikhonov et al., 1943; Tibshirani, 1996; Zou and Hastie, 2005). Along this line, the Huber loss was proposed to trade off between MAE and MSE, aiming to mitigate the issue of MSE being sensitive to outliers (Huber, 1992). All these conventional loss functions and many of their variants focus on minimizing the difference between model prediction and ground truth for each individual data sample pointwisely, but do not explicitly model and capture the relationships for multiple data samples.
@@ 2 Related Work
In this work, we focus on regression on the clean data setting without considering noises, outliers, distributional shifts, etc.
@@ 3 Method
The conventional regression task optimizes:
@@
$\mathcal{L}^{\text{MAE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}|\sigma^{f}_{\mathbf{x}_{i}}|,\quad\mathcal{L}^{\text{MSE}}_{c}=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}})^{2}.$ (2)
@@ 3.1 Motivations
Motivated by the second point, we explicitly enforce the pairwise prediction difference $f(\mathbf{x}_{i})-f(\mathbf{x}_{j})$ to be close to the pairwise label difference $y_{i}-y_{j}$ through two pairwise losses in the label space: the first directly matches the original label difference, while the second is a relaxed version with a scaling factor. Next, we prove that: the first pairwise loss is equivalent to the variance of prediction errors and the second is equivalent to negative Pearson correlation coefficient between predictions and ground truths, which addresses the concern in the first point. Moreover, the equivalent forms reduce the quadratic time complexity for the proposed pairwise losses to linear complexity. Last but not least, we provide theoretical insight between learning the pairwise label difference and learning the gradients of the ground truth function.
@@ 3.2 Pairwise Losses
When choose individual loss $\ell$ as $\frac{1}{2}$ squared error function, $\ell_{\frac{1}{2}\text{MSE}}(a,b)=\frac{1}{2}(a-b)^{2}$:
@@
$\mathcal{L}^{\text{MSE}}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}}+\bar{\sigma}^{f}_{\mathbf{x}})^{2}$
$=\frac{1}{N}\sum_{i=1}^{N}[(\sigma^{f}_{\mathbf{x}_{i}}-\bar{\sigma}^{f}_{\mathbf{x}})^{2}+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}]=\text{Var}(\sigma^{f}_{\mathbf{x}})+(\bar{\sigma}^{f}_{\mathbf{x}})^{2}.$
@@
Remark 1: The proof mainly depends on Mean Value Theorem, Limits and L’Hôpital’s rule. Besides of fitting the prediction $f(\mathbf{x}_{i})$ to be close to the ground truth $y_{i}$, that is, $f(\mathbf{x})\approx\mathcal{Y}(\mathbf{x}),~{}\forall(\mathbf{x},\mathcal{Y}(\mathbf{x}))\in\mathcal{D}$, we additionally make the model to explicitly capture the gradient (and the higher-order gradients) of the ground truth function, i.e. $\nabla^{k}f(\mathbf{x})\approx\nabla^{k}\mathcal{Y}(\mathbf{x}),~{}k=\{1,2,...\}$, with the proposed pairwise losses.
Remark 2: It should be noted that the proposed pairwise losses are designed to work with conventional pointwise regression loss (e.g. MAE), which complementarily captures pointwise label values.
@@ 4.1 Synthetic Experiments
Model: We utilize a 7-layer simple Feed Forward Neural Network (FFNN) with 5 hidden layers where neuron number for each layer is set as 100 with ELU activation function (Clevert et al., 2015).
Experimental Settings: For both of the synthetic datasets, we uniformly randomly sample half of the data as the training dataset. The full details for experimental settings are in the Appendix B.3.
@@
Table 1: Testing performance on eight different datasets/settings. Mean and standard deviation are reported. ‘Gains’ includes the percentages that GAR outperform the MAE and the best competitor (in parenthesis). The ‘p-values’ are calculated based on Student’s t-test for the ‘Gains’. p-value < 0.05 usually means significant difference. $\downarrow$ means the smaller the better; $\uparrow$ means the larger the better.
@@ 4.2 Benchmark Experiments
Experimental Settings: we follow the standard procedure for data splitting, cross validation and hyper-parameter tuning. The full details are included in Appendix B.7.
@@ 4.3 Further Analysis
We provide practical running time comparison, ablation studies on different components, sensitivity analysis on $\alpha$ and batch size for GAR and difference with data standardization in this subsection.
Running Time Comparison: we run all compared methods sequentially on an exclusive cluster node with AMD EPYC 7402 24-Core Processor 2.0 GHz. Each method runs 1500 epochs for the five tasks on the tabular datasets. The results are summarized in Table 2. From the perspective of ‘All Time’, we find that GAR is as efficient as the most efficient conventional regression loss, such as MAE, MSE, etc. From the perspective of ‘Loss Time’, we find that GAR consistently takes slightly more time than conventional regression losses (but still in the same order/level). It is consistent with our theoretical result that GAR only takes linear time complexity. On the other hand, the pairwise latent feature based regression methods, such as RankSim, ConR and RNC, are clearly slower than GAR.
Table 2: Running time (seconds/epoch) on five benchmark datasets. Repeated 1500 epochs, mean and standard deviation are reported. ‘CCS’: Concrete Compressive Strength. ’All Time’: all the time consumed per epoch. ’Loss Time’: the time for loss calculation and loss backward operation.
Dataset Concrete Compressive Strength Wine Quality Parkinson (Total) Super Conductivity IC50
Metric All Time Loss Time All Time Loss Time All Time Loss Time All Time Loss Time All Time Loss Time
MAE 0.089(0.013) 0.002(0.0) 0.169(0.03) 0.011(0.0) 0.214(0.129) 0.045(0.113) 0.587(0.038) 0.162(0.001) 0.094(0.003) 0.005(0.0)
MSE 0.09(0.014) 0.002(0.005) 0.168(0.03) 0.01(0.0) 0.21(0.024) 0.041(0.0) 0.566(0.002) 0.16(0.001) 0.095(0.003) 0.005(0.0)
Huber 0.09(0.015) 0.002(0.008) 0.168(0.03) 0.01(0.0) 0.21(0.046) 0.041(0.0) 0.581(0.038) 0.159(0.001) 0.094(0.003) 0.005(0.0)
Focal (MAE) 0.09(0.013) 0.002(0.001) 0.177(0.03) 0.016(0.0) 0.214(0.024) 0.046(0.0) 0.603(0.037) 0.178(0.001) 0.096(0.041) 0.005(0.0)
Focal (MSE) 0.089(0.012) 0.002(0.0) 0.173(0.03) 0.015(0.0) 0.216(0.046) 0.046(0.001) 0.665(0.061) 0.179(0.002) 0.094(0.003) 0.005(0.0)
RankSim 0.364(0.046) 0.275(0.045) 0.224(0.021) 0.068(0.002) 1.982(0.033) 1.81(0.03) 6.876(1.024) 6.428(0.977) 0.362(0.009) 0.271(0.006)
RNC 0.371(0.012) 0.275(0.01) 2.178(0.016) 2.014(0.015) 2.195(0.03) 2.02(0.029) 7.989(0.231) 7.545(0.168) 0.275(0.004) 0.178(0.003)
ConR 0.103(0.011) 0.014(0.004) 0.269(0.026) 0.108(0.001) 0.307(0.023) 0.138(0.001) 0.943(0.057) 0.508(0.022) 0.108(0.002) 0.02(0.0)
GAR (ours) 0.086(0.012) 0.003(0.0) 0.177(0.026) 0.026(0.0) 0.22(0.024) 0.056(0.001) 0.64(0.039) 0.213(0.002) 0.09(0.003) 0.006(0.0)
@@
Sensitivity on $\alpha$: we conduct extensive hyper-parameter search for $\alpha$ in {0.1, 0.2, 0.25, 0.4, 0.5, 0.8, 1, 1.25, 2, 2.5, 4, 5, 10}. The initial learning rate and weight decay are still searched in the same previous sets that are described in Appendix B.7. We include the full results in Appendix B.10. From the box-plot, GAR is a little sensitive to $\alpha$ and it is suggested to tune it in the [0.1, 10] range.
Sensitivity on Batch Sizes: we additionally conduct experiments on different batch sizes, {32, 64, 128, 256, 512}, for GAR on Concrete Compressive Strength, Parkinson (Total) and Parkinson (Motor) datasets. Details and results are included in Appendix B.11. GAR is also a little sensitive to batch size and suggested to be tuned in {32, 64, 128} in practice.
Difference with Data Standardization: to emphasize the difference between Eq. 53 and data centering or standardization, we include extensive discussions and experiments in the Appendix B.12 and Tab. 7. The proposed GAR enjoys similar superior performance on the datasets after data standardization as the main results in Tab. 1.
@@ 5 Limitations
The scope for this work is limited to improving learning label pairwise difference for regression on clean data setting, without considering noises, outliers, or distributional shifts. Based on the Theorem 4, we relate GAR to capture the gradient of the ground truth function. However, GAR might be misled if the training data does not reflect the underlying ground truth function. It would be an interesting future work to research the robustness for GAR under extensive settings.
@@ B.7 Experimental Setting on Real World Datasets
For tabular datasets, we uniformly randomly split 20% data as testing; the remaining 80% as training and validation, where we conduct 5-fold-cross-validation with random seed set as 123. The total training epochs is set as 100 and batch size is set as 256. The weight decay for each method is tuned in {1e-3, 1e-4, 1e-5}; we utilize SGD with momentum (set as 0.9) optimizer and tune the initial learning rate for baseline method in {1e-1, 1e-2, 1e-3, 1e-4, 1e-5}, which is stage-wised decreased by 10 folds at the end of 50-th and 75-th epoch. The switching hyper-parameter $\delta$ for Huber loss and the scaling hyper-parameter $\beta$ for Focal (MAE) or Focal (MSE) loss are tuned in {0.25, 1, 4}. The interpolation hyper-parameter $\lambda$ for RankSim is tuned in {0.5, 1, 2} and the balancing hyper-parameter $\gamma$ is fixed as 100 as suggested by their sensitivity study in their Appendix C.4 (Gong et al., 2022). The temperature hyper-parameter for RNC is tuned in {1,2,4}; the first 50 epochs are used for RNC pre-training and the remaining 50 epochs are used for fine-tuning with MAE loss. The linear combination hyper-parameters $\alpha$ is fixed as 1, $\beta$ is tuned in {0.2, 1, 4} for ConR, as suggested by the ablation studies in their Appendix A.5 (Keramati et al., 2023). The robust reconciliation hyper-parameter $\alpha$ for GAR is tuned in {0.1, 1, 10}. The model performance is evaluated with the following 4 metrics: MAE, RMSE, Pearson correlation coefficient, Spearman’s rank correlation coefficient. The best testing performance for each baseline method on each training fold is decided by its best validation performance on the same training fold over (initial learning rate, weight decay, method-special-hyper-parameter, the epoch for evaluation) combinations. For IC50 with 15 targets, we report the averaged value across all targets. The selected $\alpha$ for different datasets on different data fold are summarized in Tab. 4. For AgeDB dataset, we follow the same setting for RNC linear probe task and similar setting for training from scratch task. We include the complete details in Appendix B.8. Because RNC is a pre-training method, it doesn’t have evaluation on training from scratch task. Instead, all baselines in linear probe are based on RNC pre-trained latent features. Because RankSim and ConR are proposed to manipulate latent feature, they don’t have evaluation on linear probe task where the trainable model is linear. We also include the coefficient of determination (R2) as an evaluation metric for AgeDB in addition to the previous four evaluation metrics, to compare with prior research.
Table 4: The selected $\alpha$ on the tabular data. Standard deviations are included in the parenthesis.
Dataset Data Fold #1 #2 #3 #4 #5 Average
Concrete Compressive Strength MAE 10 10 10 10 10 10.0(0.0)
RMSE 10 1 10 10 10 8.2(3.6)
Pearson 10 1 10 10 10 8.2(3.6)
Spearman 10 1 10 10 10 8.2(3.6)
Wine Quality MAE 0.1 0.1 0.1 0.1 0.1 0.1(0.0)
RMSE 10 10 10 10 10 10.0(0.0)
Pearson 10 10 10 10 10 10.0(0.0)
Spearman 10 10 10 10 10 10.0(0.0)
Parkinson
MAE 10 1 10 1 10 6.4(4.409)
RMSE 0.1 0.1 0.1 10 0.1 2.08(3.96)
Pearson 0.1 0.1 0.1 10 0.1 2.08(3.96)
Spearman 0.1 0.1 0.1 10 0.1 2.08(3.96)
Parkinson
MAE 0.1 0.1 0.1 0.1 0.1 0.1(0.0)
RMSE 0.1 0.1 0.1 0.1 0.1 0.1(0.0)
Pearson 0.1 0.1 0.1 0.1 0.1 0.1(0.0)
Spearman 0.1 0.1 0.1 0.1 0.1 0.1(0.0)
Super
Conductivity
MAE 10 10 1 10 10 8.2(3.6)
RMSE 10 10 1 1 1 4.6(4.409)
Pearson 0.1 10 1 1 0.1 2.44(3.801)
Spearman 10 10 10 10 10 10.0(0.0)
IC50 MAE 1 1 1 1 1 1.0(0.0)
RMSE 1 1 1 0.1 0.1 0.64(0.441)
Pearson 1 1 10 10 10 6.4(4.409)
Spearman 1 1 10 10 10 6.4(4.409)